OpenStudy (anonymous):

A salmon farm has seen its sales decrease by an average of 4 percent per year from a high of 15 million pounds in 2007. During the same period, it has increased its prices by an average of 6 percent per year from $1.50 per pound in 2007.

OpenStudy (anonymous):

A. Determine how many pounds of salmon the farm sold in 2009. Then calculate the difference between dollars earned from sales in 2009 and dollars earned from sales in 2007. Round your figures to the nearest cent. I got $1.69 per pound

OpenStudy (anonymous):

B. Calculate the minimum number of pounds of salmon the farm would need to sell at 2007 prices to match its 2009 earnings. Round your answer to the nearest pound. Explain how government subsidies could help the farm become more profitable.

OpenStudy (anonymous):

I need help @Hero @dan815 @welshfella @piano_dance @misty1212 @robtobey @samanthagreer

OpenStudy (anonymous):

@abb0t @chrisdbest @Blondisaurus @Frostbite @izzyoct @oriol_00 @Mehek14 @Blondisaurus I am really desperate fan and medal

OpenStudy (anonymous):

I need help with b through d

OpenStudy (anonymous):

@Hero @dan815 @wolf1728 @WesternBurger @NathalyN @Nerdgirl

OpenStudy (anonymous):

Can you help me

OpenStudy (wolf1728):

If the 2007 amount of fish is 15,000,000 pounds, then decreasing 4% per year the 2009 amount would be 13,824,000 In part B it asks to calculate the 2009 figure but isn't that figure already determined by the wording of the problem? That's a strangely worded problem.

OpenStudy (anonymous):

I already know the pounds for 2009. I

OpenStudy (anonymous):

but I am having trouble with b

OpenStudy (anonymous):

I think it is asking how many pounds that will have to be sold in 2007 to reach the price of 2009

OpenStudy (wolf1728):

Well, that's as far as I can go. Guess you'll need someone else to help. Good luck :-)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Can ask you someone for me

OpenStudy (mathstudent55):

Part A. Sales: 2007 15,000,000 lb 2008 15,000,000 lb(0.96) = 14,400,000 lb 2009 14,400,000 lb(0.96) = 13,824,000 lb Price: 2007 $1.50/lb 2008 $1.50/lb * 1.06 = $1.59/lb 2009 $1.59/lb * 1.06 = $1.69/lb Dollars earned in 2007: 15,000,000 lb * $1.50/lb = $22,500,000 Dollars earned in 2009: 13,824,000 lb * $1.69/lb = $23,362,560 Difference: $23,362,560 - $22,500,000 = $862,560

OpenStudy (mathstudent55):

Part B. $23,362,560/(1.50/lb) = 15,575,040 lb

OpenStudy (anonymous):

Thank you so much